OpenStudy (anonymous):

A tugboat goes 160 miles upstream in 10 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current.

OpenStudy (radar):

Let x = the current speed. Let b = boat speed. Remember d=st (distance equals speed times time) 160 = 10(b-x) and 160 = 5 (b+x) now we can say: multiply 2nd equation by 2 getting 320=10b +10x now add the first equation 160 = 10b -10x 320 = 10b +10x --------------- 480 = 20b b=480/20 = 24mph (boat speed) substitute 24 for b in any of the equations above 160= 5(24 + x) 160 = 120 +5x 5x =160 - 120 5x = 40 x =8 mph (current speed)

OpenStudy (anonymous):

Thank you for showing me ALL the steps Radar.

OpenStudy (radar):

You are welcome.